OpenStudy (anonymous):

How do I find the zeros of: 3x^3 -48x

OpenStudy (anonymous):

First factor out an x

OpenStudy (anonymous):

x (3x -48)

OpenStudy (anonymous):

then the zero property tells you that either term, x or 3x^2 - 48 can equal zero to solve this, so set each equal to zero independently and solve.

OpenStudy (anonymous):

3x^2 - 48 = 0 3x ^2= 48 x^2= 16 x=4? x=0? so the zeros are x=4, and x=0?:S

OpenStudy (anonymous):

don't forget the other solution to x^2 = 16...

OpenStudy (anonymous):

x=-4 & x=4 ?

OpenStudy (anonymous):

Yes, there are three zeros.

OpenStudy (anonymous):

Well Done

OpenStudy (anonymous):

thank you!